Adv. Math. Econ. 15, 1–26 (2011)

Some various convergence results for normal integrands

Charles Castaing

Département de Mathématiques, Faculté des Sciences et Techniques Université Montpellier II, 34095 Montpellier Cedex, France (e-mail: charles.castaing@gmail.com)

Received: October 4, 2010 Revised: October 25, 2010

JEL classification: C01, C02

Mathematical Subject Classification (2010): 28B20, 60G42, 46A17, 54A20

Abstract. We prove various results for the conditional expectation of multifunctions and normal integrands. Applications to the epiconvergence of integrand reversed martingales and lower semicontinuous superadditive random sequences are presented.

Key words: conditional expectation, epiconvergence, ergodic, martingale, normal integrands, superadditive